OpenStudy (anonymous):

box contains 15 items,4 of which are defective and 11 are good. Two items are selected. What is probability that the first is good and the second defective?

OpenStudy (anonymous):

\[P(x)= ^nC_x* p^x * (1 - p)^{n - x}\] where p is the probability of success, n is the number of items, and x is the sample n=2, p=4/15 and x=1 \[P(1)= ^2C_1* (4/15)^1 * (1 - (4/15))^{2 - 1}\]\[P(1)= 2* (4/15) * (11/15))^1\]\[P(1)= (8/15) * (11/15)\]\[P(1)= 88/225\approx 0.391\]